A parallel-plate air capacitor is to store charge of magnitude 240.0 pC on each plate when the


potential difference between the plates is 42.0 V. (a) If the area of each plate is 6.80 cm2


, what is


the separation between the plates? (b) If the separation between the two plates is double the value


calculated in part (a), what potential difference is required for the capacitor to store charge of


magnitude 240.0 pC on each plate?



1
Expert's answer
2022-04-19T15:27:46-0400

"C=\frac qU=\frac{\varepsilon_0 S}d,"

a)

"d=\frac{\varepsilon_0 SU }{q}=1.05~mm,"

b)

"U'=2U=84~V."